The ratio of pharmacy technicians on duty at the hospital during for the day vs. night shifts is 32:19. 

1. Express as a percentage how many pharmacy technician staff work the night shift. Round your final answer as a percentage rounded to two decimal places.  

2. Explain in a well-written paragraph the method you’ve used to arrive at your answer. 


1
Expert's answer
2021-05-17T15:18:22-0400

1.

Percentage of pharmacy technician staff who work the night shift:

"\frac{19}{32+19}=0.3725=37.25\%"


2.

Let x is some integer, then:

total number of technicians is "32x+19x=51x"

number of night shift technicians is "19x"

So, required percentage is "19x\/51x=19\/51"